OpenStudy (anonymous):

Will Medal! Calculate | -3 + 4i |.

OpenStudy (anonymous):

-25 + 8i?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

the absolute value of a complex number is a positive real number

OpenStudy (anonymous):

\[|a+bi|=\sqrt{a^2+b^2}\] just like pythagoras ignore the \(i\) entirely

OpenStudy (anonymous):

Oh ok! So \[\sqrt{-3^{2}+ 4^{2}} = 5\]

OpenStudy (anonymous):

no minus signs

OpenStudy (anonymous):

but yeah, \(5\) is the answer

OpenStudy (anonymous):

Great! Thank you for you time and help!